# Measurements of $K_{S}^{0}$-$K_{L}^{0}$ asymmetries in the decays $\Lambda_{c}^{+}\to pK_{L,S}^{0}$, $pK_{L,S}^{0}\pi^{+}\pi^{-}$ and $pK_{L,S}^{0}\pi^{0}$

### v1

@@ Abstract
Using $e^{+}e^{-}$ annihilation data sets corresponding to an integrated luminosity of 4.5 $\text{fb}^{-1}$, collected with the BESIII detector at center-of-mass energies between 4.600 and 4.699 GeV, we report the first measurements of the absolute branching fractions $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0})=(1.67\pm 0.06\pm 0.04)\%$, $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0}\pi^{+}\pi^{-})=(1.69\pm 0.10\pm 0.05)\%$, and $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0}\pi^{0})=(2.02\pm 0.13\pm 0.05)\%$, where the first uncertainties are statistical and the second systematic. Combining with the known branching fractions of $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{0}$, we present the first measurements of the $K_{S}^{0}$-$K_{L}^{0}$ asymmetries $R(\Lambda_{c}^{+},K_{S,L}^{0}X)=\frac{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)-\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)+\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}$ in charmed baryon decays: $R(\Lambda_{c}^{+},pK_{S,L}^{0})=-0.025\pm 0.031$, $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{+}\pi^{-})=-0.027\pm 0.048$, and $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{0})=-0.015\pm 0.046$. No significant asymmetries within the uncertainties are observed.
@@ 1 Introduction
The lightest charmed baryon, $\Lambda_{c}^{+}$, provides a unique environment for studying the behavior of light di-quarks in the presence of a heavy quark Cheng:2021qpd . Its hadronic decays occur only through the weak interaction, and various theoretical models have been proposed. These include the covariant confined quark model Korner:1992wi ; Ivanov:1997ra , the pole model Cheng:1991sn ; Xu:1992vc ; Cheng:1993gf ; Xu:1992sw ; Zenczykowski:1993jm ; Sharma:1998rd , current algebra Korner:1978ec ; Uppal:1994pt , and SU(3) flavor symmetry approaches Lu:2016ogy ; Savage:1989qr ; Kohara:1991ug ; Verma:1995dk ; Chau:1995gk . Its decay falls into three categories: Cabibbo-favored (CF) decays, singly Cabibbo-suppressed decays, and doubly Cabibbo-suppressed (DCS) decays. The decay amplitudes of the CF and DCS modes are expected to be proportional to the products of the Cabibbo-Kobayashi-Maskawa elements $|V_{ud}^{*}V_{cs}|$ and $|V_{us}^{*}V_{cd}|$, respectively. The ratio of their decays is approximately of the order of $\mathcal{O}(10^{-3})$, resulting in a small branching fraction (BF) for the DCS decay and making it challenging to observe directly in experiments.
In addition to direct measurements of DCS decays, the amplitudes of DCS modes can be probed using the $K^{0}_{S}$-$K^{0}_{L}$ asymmetry in the decays into neutral kaons, which arises from the interference between CF and DCS amplitudes Bigi:1994aw ; Wang:2017gxe . The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry has been studied in the decays of charmed $D$ mesons, where the asymmetry is defined by
@@
and $X$ can be $\pi^{0}$, $\eta$, $\eta^{\prime}$, $\omega$, $\rho^{0}$ or $\phi$. A large asymmetry for $R(D^{0},K_{S,L}^{0}\pi^{0})$ was reported in a previous measurement by the CLEO experiment as $R(D^{0},K_{S,L}^{0}\pi^{0})=0.108\pm 0.025\pm 0.024$ CLEO:2007rhw , where the first uncertainty is statistical and the second is systematic. The BESIII experiment reported measurements of the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries $R(D^{0},K_{S,L}^{0}X)$, where $X=\phi$, $\eta$, $\omega$ BESIII:2022xhe . Significant asymmetries were observed in $D^{0}\to K^{0}_{L}\eta$ and $D^{0}\to K^{0}_{L}\eta^{\prime}$ decays with $R(D^{0},K_{S,L}^{0}\eta)=0.080\pm 0.022$ and $R(D^{0},K_{S,L}^{0}\eta^{\prime})=0.108\pm 0.035$, respectively. In addition, this asymmetry has been investigated for the lightest charmed strange meson, and $R(D_{s}^{+},K_{S,L}^{0}K^{+})$ was determined to be $(-2.1\pm 1.9\pm 1.6)\%$ BESIII:2019kfh . However, such measurements have not been made for the decays of charmed baryons.
@@ 2 BESIII experiment and Monte Carlo simulation
Simulated samples generated with geant4-based GEANT4:2002zbu Monte Carlo (MC) software, which includes the geometric description of the BESIII detector and the detector response performance YouZhengYun2008 ; Liang:2009zzb ; Huang:2022wuo , are used to determine detection efficiencies and to estimate potential background contributions. The simulation describes the beam energy spread and the initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc Jadach:2000ir ; Jadach:1999vf . The inclusive MC samples, corresponding to about 40 times the number of events of the data samples, include the production of $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs, open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc Jadach:2000ir ; Jadach:1999vf . The known decay modes are modeled with evtgen Lange:2001uf ; Ping:2008zz using BFs taken from the PDG ParticleDataGroup:2022pth , and the remaining unknown charmonium decays are modeled with lundcharm Chen:2000tv ; Lange:2001uf . Final state radiation from charged final state particles is incorporated using photos Richter-Was:1992hxq . For the production of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ events, the Born cross-section line shape from BESIII measurements is used BESIII:2017kqg ; BESIII:2023rwv . Exclusive $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ signal MC samples are generated with $\bar{\Lambda}_{c}^{-}$ decaying to twelve specific tag modes (as described in Section 3) and $\Lambda_{c}^{+}$ decaying to $pK^{0}_{L}$, $pK^{0}_{L}\pi^{+}\pi^{-}$ and $pK^{0}_{L}\pi^{0}$. The angular distribution of the decay $\Lambda_{c}^{+}\to pK^{0}_{L}$ is modeled with decay asymmetry parameters obtained from Ref. BESIII:2019odb . For processes from $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ channels, signal models are tunned based on the data. Additional MC samples are generated to estimate contributions from peaking background processes, where $\bar{\Lambda}_{c}^{-}$ decays into tag modes and $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $p\eta$, $pK^{0}_{S}\pi^{0}$, and $pK^{0}_{S}\pi^{+}\pi^{-}$, with $K^{0}_{S}$ and $\eta$ decaying inclusively. Each tag mode of the exclusive MC samples is generated with the same number of events.
@@ 3 Data analysis
To improve the momentum resolution, a six constraint (6C) kinematic fit is performed requiring total four-momentum conservation with respect to that of the initial $e^{+}e^{-}$ collision and constraining both masses of the tagged $\bar{\Lambda}_{c}^{-}$ and the signal $\Lambda_{c}^{+}$ to $m_{\Lambda_{c}^{+}}$. The $K^{0}_{L}$ is treated as a missing particle, and its four-momentum and mass are free in the kinematic fit. The $\chi^{2}$ of the kinematic fit for each signal mode is required to be less than the optimized value that maximizes the figure of merit $S/\sqrt{S+B}$, where $S$ and $B$ are the numbers of signal and background events from MC simulations, scaled to the data luminosity. The optimized requirements are $\chi^{2}<60$ for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\chi^{2}<25$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\chi^{2}<20$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$. The resulting $M^{2}_{\mathrm{miss}}$ distributions of the DT events are shown in Figure 1, which combine all data samples at the seven c.m. energies. Signal events are indicated by the significant peaks around the $K^{0}_{L}$ mass squared.
@@ 4 Systematic uncertainties
  • Kinematic fit. The model of the MC simulation is much simpler than the real detector performance, resulting in a difference between the data and MC simulation in the track parameters of the charged tracks BESIII:2012mpj . The helix parameters of the charged tracks are corrected, and the BFs are re-evaluated with the updated MC simulation samples. The differences from the measured BFs are taken as the systematic uncertainties associated with the kinematic fit, which are 0.5%, 1.0%, and 0.5% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
@@
  • Signal model. For $\Lambda_{c}^{+}\to pK^{0}_{L}$, the systematic uncertainty from the signal model is determined varying the decay asymmetry parameters within ±$\pm 1\sigma$. The deviation from the measured BF is found to be negligible. For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, the signal models in the nominal analysis is tunned based on the data. The possible intermediate resonances are considered in the amplitude analysis, composed of $\Sigma^{*}$, $\Delta^{*}$, $N^{*}$, $\bar{K}^{*}$ and $\rho$. The nominal amplitude models are then replaced by alternative ones with equivalent descriptions of the data. The alternative model of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ is selected by removing the insignificant intermediate resonances. For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, the amplitude fit is not stable due to the limited statistics of data. An alternative model is chosen with a similar fit quality to the nominal. The systematic uncertainties are determined to be 1.1% and 0.1% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
@@
  • Fit bias. The systematic uncertainty from the simultaneous fit is studied with 5000 sets of toy MC samples, which are simulated with all parameters from the fit model fixed. The BFs obtained from the toy samples are fitted with a Gaussian function. The deviations between the Gaussian mean value and nominal BFs are assigned as systematic uncertainties. For the decay $\Lambda_{c}^{+}\to pK_{L}\pi^{0}$, the fit bias is found to be 0.27%, while for the other two signal modes it is negligible.
Other sources of systematic uncertainties, such as the BF of $\pi^{0}\to\gamma\gamma$, are neglected due to their negligible effects. Assuming that all sources of systematic uncertainties in the BFs measurements are uncorrelated, the quadratic sums of the different sources are considered as the total systematic uncertainties, which are 2.2%, 3.1%, and 2.3% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively. Table 4 lists all the systematic uncertainties discussed above.

### v2

@@ Abstract
Using $e^{+}e^{-}$ annihilation data sets corresponding to an integrated luminosity of 4.5 $\text{fb}^{-1}$, collected with the BESIII detector at center-of-mass energies between 4.600 and 4.699 GeV, we report the first measurements of the absolute branching fractions $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0})=(1.67\pm 0.06\pm 0.04)\%$, $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0}\pi^{+}\pi^{-})=(1.69\pm 0.10\pm 0.05)\%$, and $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0}\pi^{0})=(2.02\pm 0.13\pm 0.05)\%$, where the first uncertainties are statistical and the second systematic. Combining with the known branching fractions of $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{0}$, we present the first measurements of the $K_{S}^{0}$-$K_{L}^{0}$ asymmetries $R(\Lambda_{c}^{+},K_{S,L}^{0}X)=\frac{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)-\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)+\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}$ in charmed baryon decays: $R(\Lambda_{c}^{+},pK_{S,L}^{0})=-0.025\pm 0.031$, $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{+}\pi^{-})=-0.027\pm 0.048$, and $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{0})=-0.015\pm 0.046$. No significant asymmetries with statistical significance are observed.
@@ 1 Introduction
The lightest charmed baryon, $\Lambda_{c}^{+}$, provides a unique environment for studying the behavior of light di-quarks in the presence of a heavy quark Cheng:2021qpd . Its hadronic decays occur only through the weak interaction, and various theoretical models have been proposed. These include the covariant confined quark model Korner:1992wi ; Ivanov:1997ra , the pole model Cheng:1991sn ; Xu:1992vc ; Cheng:1993gf ; Xu:1992sw ; Zenczykowski:1993jm ; Sharma:1998rd , current algebra Korner:1978ec ; Uppal:1994pt , and SU(3) flavor symmetry approaches Lu:2016ogy ; Savage:1989qr ; Kohara:1991ug ; Verma:1995dk ; Chau:1995gk . Its decays fall into three categories: Cabibbo-favored (CF) decays, singly Cabibbo-suppressed decays, and doubly Cabibbo-suppressed (DCS) decays. The decay amplitudes of the CF and DCS modes are expected to be proportional to the products of the Cabibbo-Kobayashi-Maskawa elements $|V_{ud}^{*}V_{cs}|$ and $|V_{us}^{*}V_{cd}|$, respectively. The ratio of their decays is approximately of the order of $\mathcal{O}(10^{-3})$, resulting in a small branching fraction (BF) for the DCS decay and making it challenging to observe directly in experiments.
In addition to direct measurements of DCS decays, the amplitudes of DCS modes can be probed using the $K^{0}_{S}$-$K^{0}_{L}$ asymmetry in the decays into neutral kaons, which arises from the interference between CF and DCS amplitudes Bigi:1994aw ; Wang:2017gxe . The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry has been studied in the decays of charmed $D$ mesons, where the asymmetry is defined as
@@
and $X$ can be $\pi^{0}$, $\eta$, $\eta^{\prime}$, $\omega$, $\rho^{0}$ or $\phi$. A large asymmetry of $R(D^{0},K_{S,L}^{0}\pi^{0})=0.108\pm 0.025\pm 0.024$ CLEO:2007rhw was reported in a previous measurement by the CLEO experiment, where the first uncertainty is statistical and the second is systematic. The BESIII experiment reported measurements of the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries $R(D^{0},K_{S,L}^{0}X)$, where $X=\phi$, $\eta$, $\omega$ BESIII:2022xhe . Significant asymmetries were observed in $D^{0}\to K^{0}_{L}\eta$ and $D^{0}\to K^{0}_{L}\eta^{\prime}$ decays with $R(D^{0},K_{S,L}^{0}\eta)=0.080\pm 0.022$ and $R(D^{0},K_{S,L}^{0}\eta^{\prime})=0.108\pm 0.035$, respectively. In addition, this asymmetry has been investigated for the lightest charmed strange meson, and $R(D_{s}^{+},K_{S,L}^{0}K^{+})$ was determined to be $(-2.1\pm 1.9\pm 1.6)\%$ BESIII:2019kfh . However, such measurements have not been made for the decays of charmed baryons.
@@ 2 BESIII experiment and Monte Carlo simulation
Simulated samples generated with geant4-based GEANT4:2002zbu Monte Carlo (MC) software, which includes the geometric description of the BESIII detector and the detector response performance YouZhengYun2008 ; Liang:2009zzb ; Huang:2022wuo , are used to determine detection efficiencies and to estimate potential background contributions. The simulation describes the beam energy spread and the initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc Jadach:2000ir ; Jadach:1999vf . The inclusive MC samples, corresponding to about 40 times the number of events of the data samples, include the production of $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs, open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc Jadach:2000ir ; Jadach:1999vf . The known decay modes are modeled with evtgen Lange:2001uf ; Ping:2008zz using BFs taken from the PDG ParticleDataGroup:2022pth , and the remaining unknown charmonium decays are modeled with lundcharm Chen:2000tv ; Lange:2001uf . Final state radiation from charged final state particles is incorporated using photos Richter-Was:1992hxq . For the production of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ events, the Born cross-section line shape from BESIII measurements is used BESIII:2017kqg ; BESIII:2023rwv . Exclusive $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ signal MC samples are generated with $\Lambda_{c}^{+}$ decaying to $pK^{0}_{L}$, $pK^{0}_{L}\pi^{+}\pi^{-}$ and $pK^{0}_{L}\pi^{0}$ in half of the signal events, and $\bar{\Lambda}_{c}^{-}$ decaying to the signal final states for the other half. The remaining $\bar{\Lambda}_{c}^{-}$ and $\Lambda_{c}^{+}$ are required to decay to twelve specific tag modes, as detailed in Section 3. The angular distribution of the decay $\Lambda_{c}^{+}\to pK^{0}_{L}$ is modeled with decay asymmetry parameters obtained from Ref. BESIII:2019odb . For processes from $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ channels, signal models are tunned based on the data. Additional MC samples are generated to estimate contributions from peaking background processes, where $\bar{\Lambda}_{c}^{-}$ decays into tag modes and $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $p\eta$, $pK^{0}_{S}\pi^{0}$, and $pK^{0}_{S}\pi^{+}\pi^{-}$, with $K^{0}_{S}$ and $\eta$ decaying inclusively. Each tag mode of the exclusive MC samples is generated with the same number of events.
@@ 3 Data analysis
To improve the momentum resolution, a six constraint (6C) kinematic fit is performed requiring total four-momentum conservation with respect to that of the initial $e^{+}e^{-}$ collision and constraining both masses of the tagged $\bar{\Lambda}_{c}^{-}$ and the signal $\Lambda_{c}^{+}$ to $m_{\Lambda_{c}^{+}}$. The $K^{0}_{L}$ is treated as a missing particle, and its four-momentum and mass are free in the kinematic fit. The $\chi^{2}$ of the kinematic fit for each signal mode is required to be less than the optimized value that maximizes the figure of merit $S/\sqrt{S+B}$, where $S$ and $B$ are the numbers of signal and background events from the inclusive MC samples, scaled to the data luminosity. Here, the BF of the signal modes are assumed to be the same as the measured ones of $\Lambda_{c}^{+}\to K^{0}_{S}X$. The optimized requirements are $\chi^{2}<60$ for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\chi^{2}<25$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\chi^{2}<20$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$. The resulting $M^{2}_{\mathrm{miss}}$ distributions of the DT events are shown in Figure 1, which combine all data samples at the seven c.m. energies. Signal events are indicated by the significant peaks around the $K^{0}_{L}$ mass squared.
@@ 4 Systematic uncertainties
  • Kinematic fit. The model of the MC simulation is much simpler than the real detector performance, resulting in a difference between the data and MC simulation in the track parameters of the charged tracks BESIII:2012mpj . The correction parameters are obtained through a data-driven method using control samples of $e^{+}e^{-}\to p\bar{p}\pi^{+}\pi^{-}$ and $e^{+}e^{-}\to K^{+}K^{-}\pi^{+}\pi^{-}$. The helix parameters of the charged tracks are corrected, and the BFs are re-evaluated with the updated MC simulation samples. The differences from the measured BFs are taken as the systematic uncertainties associated with the kinematic fit, which are 0.5%, 1.0%, and 0.5% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
@@
  • Signal model. For $\Lambda_{c}^{+}\to pK^{0}_{L}$, the systematic uncertainty from the signal model is determined varying the decay asymmetry parameters within ±$\pm 1\sigma$. The deviation from the measured BF is found to be negligible. For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, the signal models in the nominal analysis is tunned based on the data. The possible intermediate resonances are considered in the amplitude analysis, composed of $\Sigma^{*}$, $\Delta^{*}$, $N^{*}$, $\bar{K}^{*}$ and $\rho$. The nominal amplitude models are then replaced by alternative ones with equivalent descriptions of the data. The alternative model of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ is selected by including other additional insignificant resonance. In the case of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, the amplitude fit is not stable due to the limited statistics of data. The same components of resonances are employed as in the nominal case, with a different fit solution that yields a negative log-likelihood value agreeing closely with the nominal one. The systematic uncertainties are determined to be 1.1% and 0.9% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
@@
  • Fit bias. The systematic uncertainty from the simultaneous fit is studied with 5000 sets of toy MC samples, which are simulated with all parameters from the fit model fixed. The BFs obtained from the toy samples are fitted with a Gaussian function. The deviations between the Gaussian mean value and nominal BFs are assigned as systematic uncertainties. For the decay $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, the fit bias is found to be 0.3%, while for the other two signal modes it is negligible.
Other sources of systematic uncertainties, such as the BF of $\pi^{0}\to\gamma\gamma$, are neglected due to their negligible effects. Assuming that all sources of systematic uncertainties in the BFs measurements are uncorrelated, the quadratic sums of the different sources are considered as the total systematic uncertainties, which are 2.2%, 3.1%, and 2.5% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively. Table 4 lists all the systematic uncertainties discussed above.